OpenStudy (anonymous):

You find a fossilized dinosaur egg. It is spherical with a diameter of 17 cm. What is the volume of the egg?

jigglypuff314 (jigglypuff314):

volume of a sphere: (4/3)pi (r^3) radius = 1/2 diameter so when diameter = 17 then radius = 8.5 so solve --> (4/3) pi ( (8.5)^3 ) =

OpenStudy (anonymous):

so is it 2572.44 ?

jigglypuff314 (jigglypuff314):

that's what I got :)

OpenStudy (anonymous):

Awesome! what about this one? You find a fossilized animal footprint in Montana. The footprint is round and has a diameter of 42 cm. What is the area of the footprint?

jigglypuff314 (jigglypuff314):

diameter means that the over all shape is a circle area of circle = pi r^2 radius = 1/2 diameter so when diameter = 42 then radius = 21 so solve --> pi (21)^2 =

OpenStudy (anonymous):

1385.44?

jigglypuff314 (jigglypuff314):

that's what I got :)

OpenStudy (anonymous):

A paleontologist finds a fossilized reptile scale. It is triangular in shape with a base of 1.5 cm and a height of 2.2 cm. What is the area of the scale?

jigglypuff314 (jigglypuff314):

area of triangle = (1/2)(base)(height) so with what you're given --> (1/2)(1.5)(2.2) =

OpenStudy (anonymous):

1.65?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

A fossilized animal footprint has three rectangular toes. Toe A has a length of 9 cm and a width of 6.5 cm. Toe B has a length of 8.7 cm and a width of 6 cm. Toe C has a length of 9.4 cm and a width of 7.2 cm. What is the surface area of the bottom of each toe?

OpenStudy (anonymous):

I have no idea where they get these scenarios..

jigglypuff314 (jigglypuff314):

area of rectangle = (length)(width) Toe A = (9)(6.5) = Toe B = (8.7)(6) = Toe C = (9.4)(7.2) =

OpenStudy (anonymous):

a = 58.5 b = 52.2 c = 67.68 The teeth of a newly discovered ancient animal are shaped like cylinders. A paleontologist finds that Tooth A is 6 cm long with a radius of 3 cm. Tooth B is 5.5 cm long with a radius of 3.5 cm. Which tooth has a larger surface area?

jigglypuff314 (jigglypuff314):

correct :) and now... surface area of cylinder = 2 pi r^2 + 2h pi r so Tooth A = 2 pi (3^2) + 2(6)pi(3) = Tooth B = 2pi(3.5^2) + 2(5.5)pi(3.5) = then compare the two...

OpenStudy (anonymous):

a = 169.65 b = 197.92 ?

jigglypuff314 (jigglypuff314):

so B is greater than A therefore B would be the answer they want :)